# Adaptive-TMLE for the Average Treatment Effect based on Randomized Controlled Trial Augmented with Real-World Data

### v1

@@ 8.1 Augmenting both the treatment and control arm
Figure 3: Comparing the MSE, relative MSE, 95% confidence interval coverage of A-TMLE, ES-CVTMLE, and TMLE estimator for increasing sample sizes of both the RCT and external data. The reference (red dashed line) in the relative MSE plots are the MSEs of the RCT-only estimator. The plots in the right-most column show the average proportion of cross-validation folds the ES-CVTMLE estimator pooled the external data.
Figure 3 shows that in scenarios with complex bias, ES-CVTMLE fails to achieve any efficiency gain, and at times, it underperforms by having a higher MSE compared to that of a TMLE on the RCT data alone. This may stem from not estimating the bias accurately or as a result of its cross-validation method not fully leveraging the data due to the requirements for bias and variance estimation for the selection criteria. Conversely, A-TMLE continues to secure a notable efficiency gain while maintaining valid type 1 error control. Table 2 shows the average width of confidence interval lengths across varying sample sizes and simulation runs of A-TMLE as a percentage of that of other methods. In scenarios (a) and (b), the confidence intervals produced by A-TMLE are much smaller. For scenarios (c) and (d), the reduction in widths is smaller, but still superior compared with competitors.
Table 2: A-TMLE’s 95% CI width as a percentage of other methods’.
Method Scenario (a) Scenario (b) Scenario (c) Scenario (d)
ES-CVTMLE 66.1% 60.4% 98.3% 98.2%
Regular TMLE 59.1% 61.2% 99.1% 99.1%
PROCOVA 59.0% 61.1% 99.0% 99.0%
RCT-only 59.0% 61.1% 99.0% 99.0%
Figure 4: Comparing the MSE, relative MSE, 95% confidence interval coverage of A-TMLE, ES-CVTMLE, and TMLE estimator for increasing sample sizes of both the RCT and external data. The reference (dashed red line) in the relative MSE plots are the MSEs of the RCT-only estimator. The plots in the right-most column show the average proportion of cross-validation folds the ES-CVTMLE estimator pooled the external data.
8.2 Augmenting using only external control subjects
In the second simulation study, we focus on a scenario with a rare binary outcome, augmenting only external controls. The incidence rate of this outcome ranges from 5% to 8% in the pooled data. Details of the data-generating process are provided in Appendix B. The results, displayed in Figure 4, indicate that A-TMLE remains the optimal choice.
@@ 9 Discussion
In this article, we introduced adaptive-TMLE for the estimation of the average treatment effect using combined data from randomized controlled trials and real-world data. The target parameter is conservatively designed to fully respect the RCT as the gold standard for treatment effect estimation without requiring additional identification assumptions beyond those inherent to RCTs. The only extra requirement is that every individual in the combined dataset must have a non-zero probability of trial enrollment, a condition typically ensured during the sampling phase for external patients. Despite focusing on a conservative target parameter where an efficient estimator might still lack power, our proposed A-TMLE demonstrated potential efficiency gains. These gains are primarily driven by the complexity of the bias working model rather than its magnitude, which is crucial as the bias might be simple yet substantial. The working model for the bias and pooled-ATE estimand are data-adaptively learned, which may be much smaller than the nonparametric model, thus allowing efficiency gain. Importantly, since we do not impose extra assumptions, virtually any external data, even those with a different outcome from the RCT, could be utilized. This approach could extend to scenarios involving surrogate outcomes, where the difference in the outcomes might be adequately approximated as a function of covariates and treatment. Additionally, A-TMLE could be useful when researchers have access to a well-designed observational study for identifying the desired causal effect, alongside other observational data potentially subject to unmeasured confounding. In these cases, combining data sources could yield a more accurate estimator of the causal effect. Extending this approach to handle multiple external data sources could be achieved either by stratifying based on the study to learn a separate bias function for each, or by pooling across studies to jointly learn a bias function. One limitation of A-TMLE is that, although theoretically the efficiency gain is driven by the complexity of the bias working model, it remains unclear in practice how much efficiency gain one should expect. For instance, our simulations have shown situations where A-TMLE is almost twice as efficient, yet other cases where the efficiency gain is small, although still better than alternative approaches. Future research should explore how much efficiency gain one can expect in finite samples. Another limitation is the challenge of determining how to enforce a minimal working model without compromising efficiency gains. Without such a minimal working model, one runs the risk of introducing bias if cross-validation overly favors a sparse model when the true model is actually complex. A better understanding of this trade-off is necessary to provide concrete guidelines on the minimal working model.
@@ Acknowledgement
The authors thank members of the Joint Initiative for Causal Inference (JICI) working group for insightful comments and discussions. This work was funded by Novo Nordisk.
@@
First consider the case that we do not assume $S\perp W$ but possibly have a model on $g(A\mid S,W)$. Then the efficiency bound corresponds with the one in the nonparametric model due to the target parameter only depending on $P_{W}$ and $\bar{Q}$, and that the tangent space of $g$ is orthogonal to the tangent space of these nuisance parameters. Therefore, it suffices to derive the influence curve of the empirical plug-in estimator acting as if $W$ is discrete, which then also establishes the general efficient influence curve by an approximation argument. The influence curve of the empirical plug-in estimator $\Psi(P_{n})$ is straightforward to derive by using that the canonical gradient/influence curve of the empirical estimate of $E(Y\mid W=w,A=a,S=s)$ is given by $E(Y\mid W=w,A=a,S=1)=I(S=1,A=a,W=w)/p(S=1,W=w,A=a)(Y-\bar{Q}(S,W,A))$, while we use the empirical measure of $W_{1},\ldots,W_{n}$ as estimator of $P_{W}$. Consider now the statistical model that also assumes that $S\perp W$. The only way this model assumption can affect the estimator is that it might yield a different estimator of $P_{W}$ than the empirical measure. However, the empirical distribution of $W$ is still an efficient estimator of $P_{W}$, even when $S$ is independent of $W$. In other words, if we have two samples of $W$ from the same population, then the efficient estimator of the marginal distribution of $W$ is still the empirical measure of the combined sample on $W$. The proof for $\Psi_{2}$ is similar. If we assume $S\perp W$, then the canonical gradient for $\Psi_{2}$ changes a little due to an efficient estimator of $P_{W\mid S=1}$ should now use the empirical measure of $W$ for the combined sample, instead of using the empirical measure of $W$, given $S=1$. ∎
@@
The proof is analogous to the proof of Lemma 6 by replacing $A$ with $S$. Specifically, in Lemma 6, we show the loss function for learning the working model for the conditional effect of $A$ on $Y$. Here, in Lemma 3, we have a similar loss function for learning the conditional effect of $S$ on $Y$. ∎
@@
Proof of lemma 5.
The $p_{Y\mid S,W,A}$-score component of $D_{{\cal M}_{w,2},P}^{\#}$ is given by
$D_{{\cal M}_{w,2},\beta,P}^{\#}$$\equiv\left\{E_{P}\Pi_{P}(0\mid W,0)\frac{d}{d\beta(P)}\tau_{w,\beta(P)}(W,0)\right\}^{\top}D_{\beta,P}(O)$
$-\left\{E_{P}\Pi_{P}(0\mid W,1)\frac{d}{d\beta(P)}\tau_{w,\beta(P)}(W,1)\right\}^{\top}D_{\beta,P}(O)$
The $p_{W}$-score component of $D_{{\cal M}_{w,2},W,P}^{\#}$ is given by
$D^{\#}_{{\cal M}_{w,2},W,P}\equiv\Pi_{P}(W,0)\tau_{w,\beta(P)}(W,0)-\Pi_{P}(W,1)\tau_{w,\beta(P)}(W,1)-\Psi_{{\cal M}_{w,2}}^{\#}(P).$
Finally we need the contribution from the dependence of $\Psi^{\#}_{{\cal M}_{w,2}}(P)$ on the conditional distribution $\Pi_{P}$. The influence curve of $\Pi_{P}(0\mid w,a)$ is given by $D_{{\cal M}_{w},\Pi_{P},(w,a),P}=I(W=w,A=a)/P(w,a)(I(S=0)-\Pi_{P}(0\mid w,a))$. So $\Pi_{P}$-score component of $D_{{\cal M}_{w,2},P}^{\#}$ is given by
[…]
=$=\left\{\frac{I(A=0)}{g_{P}(0\mid W)}\tau_{w,\beta(P)}(W,0)-\frac{I(A=1)}{g_{P}(1\mid W)}\tau_{w,\beta(P)}(W,1)\right\}(I(S=0)-\Pi_{P}(0\mid W,A)).$
Now note that $I(S=0)-\Pi(0\mid W,A)=-(I(S=1)-\Pi(1\mid W,A))$. Therefore, we have
$D^{\#}_{{\cal M}_{w},\Pi_{P},P}=\left\{\frac{A}{g_{P}(1\mid W)}\tau_{w,\beta(P)}(W,0)-\frac{1-A}{g_{P}(0\mid W)}\tau_{w,\beta(P)}(W,1)\right\}(S-\Pi_{P}(1\mid W,A)).$
So, we have every component in
$D^{\#}_{{\cal M}_{w,2},P}=D^{\#}_{{\cal M}_{w,2},W,P}+D^{\#}_{{\cal M}_{w,2},\Pi,P}+D^{\#}_{{\cal M}_{w,2},\beta,P}.$
Proof of lemma 6.
We have $\bar{Q}^{r}_{P}=\theta(P)+AT_{P}$, where $\theta(P)=E_{P}(Y\mid A=0,W)$ and $T_{P}=E_{P}(Y\mid A=1,W)-E_{P}(Y\mid A=0,W)$. We have to compute a projection of $\bar{Q}^{r}_{P}$ onto the linear space $\theta(W)+A\sum_{j}\beta(j)\phi_{j}$. We can write
$\theta(W)+A\sum_{j}\beta(j)\phi_{j}=\left(\theta(W)+g(W)\sum_{j}\beta_{j}\phi_{j}\right)+(A-g(W))\sum_{j}\beta(j)\phi_{j}.$
Thus, the linear space on which we are projecting is an orthogonal sum of $L^{2}(P_{W})$ and the linear span of $\{(A-g(W))\phi_{j}:j\}$. In $L^{2}(P)$, functions of $W$ are orthogonal to the linear span of $\{(A-g(W))\phi_{j}:j\}$. Therefore the projection of $\bar{Q}_{P}$ on this orthogonal sum space is given by $E(\bar{Q}_{P}\mid W)=E_{P}(Y\mid W)$ plus the projection of $\bar{Q}_{P}=\theta(P)+AT_{P}$ onto $\{(A-g(W))\phi_{j}:j\}$. The projection of $\theta(P)$ on this linear span equals zero since a function of $W$ is orthogonal to $(A-g(W))\phi_{j}$. The projection of $AT_{P}$ is given by $(A-g(W))\sum_{j}\beta_{P}(j)\phi_{j}$ with $\beta(P)=\arg\min_{\beta}E_{P}(AT_{P}(W)-(A-g(W))\sum_{j}\beta_{j}\phi_{j}(W))^{2}$. We can write $AT_{P}=(A-g(W))T_{P}(W)+g(W)T_{P}(W)$. So $\beta(P)=\arg\min_{\beta}E_{P}(A-g(W))^{2}(T_{P}(W)-\sum_{j}\beta_{j}\phi_{j}(W))^{2}$, which equals $\arg\min_{\beta}E_{P}g_{P}(1-g_{P})(W)(T_{P}(W)-\sum_{j}\beta_{j}\phi_{j}(W))^{2}$. This proves the lemma. ∎
Proof of lemma 7.
We want to derive the canonical gradient of $\beta(P)$ defined on a nonparametric model. Our starting point is that
$\beta(P)=\arg\min_{\beta}E_{P}\left(Y-(A-g_{P}(W))\sum_{j}\beta_{j}\phi_{j}(W)\right)^{2}.$
@@
$U(P,g_{P},\beta)=E_{P}(A-g_{P}(W)){\bf\phi}(W)\left(Y-(A-g_{P}(W))\sum_{j}\beta(j)\phi_{j}\right)=0.$
@@
where $D_{1,P}$ is the canonical gradient of $P\rightarrow U(\beta_{1},P,g_{1})$ at $\beta_{1}=\beta(P)$ and $g_{1}=g_{P}$, and $D_{2,P}$ is the canonical gradient of $P\rightarrow U(\beta_{1},P_{1},g_{P})$ at $\beta_{1}=\beta(P)$ and $P_{1}=P$. Since $U(\beta_{1},P,g_{1})$ is just a mean parameter with respect to $P$ (like $E_{P}D$ which has canonical gradient $D-E_{P}D$), it follows that
$D_{1,P}=(A-g_{P}(W)){\bf\phi}(W)\left(Y-(A-g_{P}(W))\sum_{j}\beta_{P}(j)\phi_{j}(W)\right).$
@@
$U(\beta_{1},P_{1},g_{P})=E_{P_{1}}(A-g_{P}(W)){\bf\phi}(W)\left(\bar{Q}_{P_{1}}-(A-g_{P}(W))\sum_{j}\beta_{P_{1}}(j)\phi_{j}\right).$
@@
$=\int_{w,a}p_{1}(w,a)(a-g_{P}(w)){\bf\phi}(w)\bar{Q}_{1}(w,a)-\int_{w,a}p_{1}(w,a)(a-g_{P}(w))^{2}{\bf\phi}(w)\sum_{j}\beta_{1}(j)\phi_{j}(w)$
at $p_{1}=p$. We know that the canonical gradient of $P\rightarrow g_{P}(w)$ is given by $I(W=w)/p(w)(A-g_{P}(W))$. So, by the delta-method, the canonical gradient is given by
@@
$D_{1,P}+D_{2,P}$ $=(A-g_{P}(W)){\bf\phi}(W)\left(Y-(A-g_{P}(W))\sum_{j}\beta_{P}(j)\phi_{j}(W)\right)-(A-g_{P}(W)){\bf\phi}(W)\tilde{\theta}_{P}(W)$
$=(A-g_{P}(W)){\bf\phi}(W)\left(Y-\tilde{\theta}_{P}(W)-(A-g_{P}(W))\sum_{j}\beta_{P}(j)\phi_{j}(W)\right).$

### v2

@@ 8 Simulation Studies
Figure 2: Comparing the MSE, relative MSE, 95% confidence interval coverage of A-TMLE, ES-CVTMLE, and TMLE estimator for increasing sample sizes of both the RCT and external data. The reference (red dashed line) in the relative MSE plots are the MSEs of the RCT-only estimator. The plots in the right-most column show the average proportion of cross-validation folds the ES-CVTMLE estimator pooled the external data. Scenarios (c) and (d) represent two different data-generating-processes where the conditional average RCT-enrollment effect involves indicator functions, higher-order polynomials, and higher-way interaction terms.
In some practical settings, the bias function may be highly non-linear and may not be well-approximated by a linear combination of main terms. Figure 2 shows results for such scenarios. In particular, the true bias terms in scenarios (c) and (d) involve indicator functions, higher-order polynomials, and higher-way interactions. The data generating processes are provided in Appendix B. The use of more flexible nonparametric learning algorithms for estimating the bias function is necessary in these settings. We utilize HAL with a rich set of first-order spline basis functions to data-adaptively learn the bias working model. In scenarios with complex bias, ES-CVTMLE fails to achieve any efficiency gain, and at times, it underperforms by having a higher MSE compared to that of a regular TMLE on the RCT data alone. This may stem from not estimating the bias accurately or as a result of its cross-validation method not fully leveraging the data due to the requirements for bias and variance estimation for the selection criteria. On the other hand, A-TMLE continues to offer efficiency gain while having close to 95% coverage. However, the efficiency gain is less compared to scenarios (a) and (b) with simpler bias functions, which is expected given the increased complexity of the bias.
8.1 Practical violations of trial enrollment positivity
The causal interpretation of the target parameter $\Psi$ requires that the external population either coincides with or is a subset of the trial-eligible population [dahabreh_extending_2019]. In Section 2, we further emphasized that to achieve maximal efficiency in estimating the target parameter $\Psi$, the dependence of trial enrollment on patient baseline covariates should be made as weak as possible by design when selecting external patients. However, the positivity assumption (A3), which requires that every patient has a non-zero probability of being enrolled in the trial, may still be violated practically. To evaluate the performance of A-TMLE under such conditions, we designed a simulation setup with trial enrollment near positivity violations, making the covariate distributions between the RCT and external data differ significantly. Specifically, we considered two scenarios where the smallest trial enrollment probability $P_{0}(S=1\mid W)$ was approximately 0.05 and 0.003, respectively.
Figure 3: Settings with trial enrollment positivity violations. The two subplots on the left display the densities of trial enrollment probabilities. The parameter $\gamma$ is defined as the infimum of the trial enrollment probability, $P_{0}(S=1\mid W)$.
As shown in Figure 3, A-TMLE remains to achieve the largest efficiency gain with the smallest MSE compared to other estimators. This is likely due to its avoidance of inverse weighting by the trial enrollment probability, as discussed in Section 5, making it relatively more stable in near positivity violation settings. Interestingly, in this simulation setting, ES-CVTMLE never selects the pooled-ATE estimand but still achieves a moderate efficiency gain compared to an efficient estimator applied to the RCT data alone. This occurs because when ES-CVTMLE rejects external data, it estimates the parameter $\Psi$, which quantifies the treatment effect in the trial while averaging over the covariate distribution in the pooled data. In this scenario, because certain patient subgroups underrepresented in the trial are more prevalent in the external data, such increased representation allows ES-CVTMLE to achieve moderate efficiency gain by averaging over a more representative sample of the target population.
9 Data application
We applied our method to augment the DEVOTE trial with external data from Optum’s de-identified Clinformatics database. The DEVOTE trial is a double-blind, treat-to-target, event-driven study designed to evaluate the cardiovascular safety of degludec, an ultralong-acting, once-daily basal insulin approved by the FDA for various patient populations with diabetes [marso_efficacy_2017]. Its primary objective was to determine whether degludec posed no greater cardiovascular risk than glargine, using a pre-specified noninferiority hazard ratio margin of 1.3 (allowing up to 30% excess risk). The treatment arm ($n=3,818$) included patients receiving insulin degludec, while the comparator arm ($n=3,817$) consisted of patients receiving insulin glargine. The primary endpoint was a composite outcome for major adverse cardiac events (MACE), which included death from cardiovascular causes, nonfatal myocardial infarction, and nonfatal stroke. Our scientific question of interest focuses on the 365-, 540-, and 730-day (1-, 1.5-, 2-year) risk differences in MACE between the treatment and comparator arms.
Augmenting the trial with external data to improve efficiency in this context is especially compelling for several reasons. First, there is a hypothesis that insulin degludec may reduce the risk of MACE compared to insulin glargine. However, achieving sufficient statistical power to test for superiority often requires much larger cohorts and longer follow-up periods, typically three to five years, to observe relatively rare events such as MACE [regier_more_2016]. Second, once drugs like insulin degludec meet the pre-specified noninferiority margin and receive FDA approval, substantial amounts of real-world data often become available. Such data may provide evidence of cardiovascular safety advantages of the given drug over its alternatives. In light of these findings, continuing to keep patients in the comparator arm without allowing treatment switches to a safer option might raise ethical concerns, particularly when the sole purpose for adhering to the assigned treatment status is to achieve sufficient power for a superiority test [lin_propensity_2018]. These considerations highlight the potential benefits of integrating external data with RCTs. By augmenting the DEVOTE trial with real-world data, researchers may enhance statistical power, explore opportunities for establishing treatment superiority, and improve the generalizability of the trial findings [dahabreh_extending_2019].
Challenges in augmenting the DEVOTE trial with external data include inconsistencies in outcome measurement across data sources, non-concurrency, potential unmeasured confounding in the external data, and other differences between trial settings and routine care. For example, in the DEVOTE trial, adverse events were closely monitored and confirmed by an event-adjudication committee, whereas in the external data, some events may have been missed or inconsistently recorded. The trial concluded in 2017, but we extended the external data sampling timeframe to 2021 to ensure a decent number of patients receiving insulin degludec in the external cohort. To establish comparable cohorts between the RCT and observational data we emulate the DEVOTE trial following the framework detailed in [hernan_using_2016]. We translated and applied the trials inclusion and exclusion criteria and used the initiation of insulin degludec or insulin glargine as the study start date. Although efforts have been made to address potential sources of bias, concerns about potential residual heterogeneity still remain, such as systematic differences in adherence between the controlled trial setting and routine clinical care. These differences can lead to non-comparable exposure definitions, violating the consistency assumption, which further underscores the necessity of applying our method that does not rely on assumptions beyond those that are known to likely hold true based on the design of the RCT. Censoring events were handled using the inverse probability of censoring weights technique documented in Appendix B.
Figure 4: Estimated 365-, 540-, and 730-day risk differences (%) of MACE between the insulin degludec and insulin glargine groups, with corresponding 95% confidence intervals. ‘RCT’ refers to results from an augmented inverse propensity score weighted estimator applied to DEVOTE trial data only. ‘Pooled’ refers to results from the same estimator applied to combined data from the DEVOTE trial and external Optum data. ‘ES-CVTMLE’ and ‘A-TMLE’ refer to results from the ES-CVTMLE and A-TMLE estimator applied to the pooled data, respectively.
Estimated risk difference (%) of MACE with 95% CIs
Estimator Study design 365-day 540-day 730-day
AIPW DEVOTE -0.190 (-1.177, 0.798) -0.704 (-1.888, 0.481) -1.809 (-4.065, 0.447)
AIPW DEVOTE + Optum -0.906 (-1.565, -0.248) -1.031 (-1.834, -0.227) 0.396 (-0.705, 1.497)
ES-CVTMLE DEVOTE + Optum -0.148 (-1.137, 0.842) -0.697 (-1.883, 0.488) -1.594 (-3.692, 0.718)
A-TMLE DEVOTE + Optum -0.586 (-1.118, -0.054) -0.645 (-1.288, -0.001) -1.804 (-3.855, 0.246)
Table 3: Estimated risk differences of MACE between the insulin degludec and insulin glargine groups, with corresponding 95% confidence intervals. Note that the AIPW estimator applied to the pooled data may be biased. denotes statistical significance under $\alpha=0.05$ for a two-sided test.
Figure 4 shows the estimated 1-, 1.5-, and 2-year risk differences of MACE using different estimation strategies, with corresponding values presented in Table 3. Specifically, the label ‘RCT’ refers to estimates obtained using augmented inverse propensity weighted (AIPW) estimators applied to the RCT data. The label ‘Pooled’ corresponds to AIPW estimates of pooled-ATE estimand $\tilde{\Psi}$, which, as discussed in Section 2, may be biased had the mean exchangeability over studies assumption been violated. In fact, we do observe some evidence suggesting that this might be the case, since the ‘Pooled’ effect estimate shows the opposite direction to the RCT estimate for 730-day. Even though the confidence interval overlaps between the two, large deviations from the RCT only estimate are still concerning because in this case, we expect RCT to give unbiased estimate, so the point estimates should be close. This further highlights the need to use estimators including ES-CVTMLE and A-TMLE which do not rely on those assumptions. As one could see, the ES-CVTMLE and A-TMLE point estimates are more aligned with the RCT estimate. Comparing the two, A-TMLE consistently produces narrower confidence intervals. For the 1- and 1.5-year risk estimates, A-TMLE provides statistically significant results supporting the superiority of insulin degludec in reducing MACE risk. The larger confidence intervals at the 2-year mark may be attributed to a higher number of censored patients, leading to increased variability in the estimates. We also provide the values for the point estimates and corresponding 95% confidence intervals in Table 3.
@@ 10 Discussion
In this article, we proposed an estimator for the average treatment effect using RCT data combined with RWD. We considered a target parameter that fully respects the RCT as the gold standard for treatment effect estimation without relying on extra identification assumptions beyond those that are known to be true for RCTs. The only additional assumption is that every individual in the pooled population must have a non-zero probability of participating in the RCT, a condition typically could be made more plausible during the sampling of external real-world patients. Despite focusing on a conservative target parameter for which an efficient estimator might still lack statistical power, our proposed A-TMLE demonstrated potential efficiency gains. These gains are primarily driven by the complexity of the bias working model rather than the bias magnitude, which is a desirable property, particularly in scenarios where the bias is expected to be large.
From a theoretical perspective, although A-TMLE gave up some regularity along perturbation paths of the data distribution that fall outside the learned working model, the oracle model, defined as the model that contains the true data distribution $P_{0}$, includes only relevant perturbations of the data distribution, so such loss of regularity is generally not a problem. Nonetheless, since the data-adaptive working model approximates the oracle model, going for a larger working model than what cross-validation suggests might result in finite sample bias reductions. Therefore, in practice, one may consider strategies including undersmoothing [vdl_efficient_2023] or enforcing a minimal working model, such as a main-term only minimal working model, to further protect against finite sample bias. In this manner, the family of A-TMLEs is rich enough to determine a most sensible trade-off between regularity (i.e., possible bias) and efficiency gain.
We highlight some possible extensions of A-TMLE. First, since we do not impose extra assumptions, virtually any external data, even those with an outcome defined slightly differently from the RCT, could be utilized. This approach could extend to scenarios involving surrogate outcomes, where the difference in the outcomes might be adequately approximated as a function of covariates and treatment. Additionally, A-TMLE could be useful when researchers wish to augment a well-designed observational study, where the desired causal effect could be identified, with other observational data potentially subject to unmeasured confounding. Extending this approach to handle multiple external data sources could be achieved either by stratifying based on the choice of external study to learn a separate bias function for each external data source, or by pooling across the external studies to jointly learn a bias function.
One limitation of A-TMLE is that, although theoretically the efficiency gain is driven by the complexity of the bias working model, it remains unclear in practice how much efficiency gain one should expect. For instance, our simulations have shown situations where A-TMLE is almost twice as efficient, yet other cases where the efficiency gain is small. Future research could focus on characterizing the expected efficiency gain in finite samples and identifying conditions under which substantial gains are most likely. Another challenge is determining how to enforce a minimal working model without compromising efficiency gains. Generally, adding more terms to the working model reduces bias but may limit potential efficiency gain. However, without such a minimal working model, one runs the risk of introducing bias if cross-validation overly favors a sparse model when the true model is actually complex. Addressing this may involve implementing undersmoothing methods that find the right trade-off between finite sample bias and efficiency. Even with robust estimation mechanisms in place, conducting a sensitivity analysis remains important to provide an additional layer of confidence. Future work could explore the possibility of integrating, for example, negative control outcomes (or treatments), as part of A-TMLE.
@@ Acknowledgement
The authors thank members of the Joint Initiative for Causal Inference (JICI) working group for insightful comments and discussions. We also extend our gratitude to the two anonymous reviewers for their valuable suggestions, which greatly improved the manuscript. This work was funded by Novo Nordisk.
[…]
@@
First consider the case that we do not assume $S\perp W$ but possibly have a model on $g(A\mid S,W)$. Then the efficiency bound corresponds to the one in the nonparametric model due to the target parameter only depending on $P_{W}$ and $Q$, and that the tangent space of $g$ is orthogonal to the tangent space of these nuisance parameters. Therefore, it suffices to derive the influence curve of the empirical plug-in estimator acting as if $W$ is discrete, which then also establishes the general efficient influence curve by an approximation argument. The influence curve of the empirical plug-in estimator $\Psi(P_{n})$ is straightforward to derive by using that the canonical gradient/influence curve of the empirical estimate of $E(Y\mid W=w,A=a,S=s)$ is given by $E(Y\mid W=w,A=a,S=1)=I(S=1,A=a,W=w)/p(S=1,W=w,A=a)(Y-Q(S,W,A))$, while we use the empirical measure of $W_{1},\ldots,W_{n}$ as estimator of $P_{W}$. Consider now the statistical model that also assumes that $S\perp W$. The only way this model assumption can affect the estimator is that it might yield a different estimator of $P_{W}$ than the empirical measure. However, the empirical distribution of $W$ is still an efficient estimator of $P_{W}$, even when $S$ is independent of $W$. In other words, if we have two samples of $W$ from the same population, then the efficient estimator of the marginal distribution of $W$ is still the empirical measure of the combined sample on $W$. The proof for $\Psi_{2}$ is similar. If we assume $S\perp W$, then the canonical gradient for $\Psi_{2}$ changes a little due to an efficient estimator of $P_{W\mid S=1}$ should now use the empirical measure of $W$ for the combined sample, instead of using the empirical measure of $W$, given $S=1$. ∎
@@
We have $\bar{Q}_{P}(W,A)=\tilde{\theta}_{P}(W)+A\tau_{A,P}(W)$, where $\tilde{\theta}_{P}(W)=E_{P}(Y\mid A=0,W)$ and $\tau_{A,P}(W)=E_{P}(Y\mid A=1,W)-E_{P}(Y\mid A=0,W)$. We have to compute a projection of $\bar{Q}_{P}$ onto the linear space $\tilde{\theta}(W)+A\sum_{j}\beta(j)\phi_{j}(W)$. We can write
$\tilde{\theta}(W)+A\sum_{j}\beta(j)\phi_{j}(W)=\left(\tilde{\theta}(W)+g(1\mid W)\sum_{j}\beta(j)\phi_{j}(W)\right)+(A-g(1\mid W))\sum_{j}\beta(j)\phi_{j}(W).$
Thus, the linear space on which we are projecting is an orthogonal sum of $L^{2}(P_{W})$ and the linear span of $\{(A-g(1\mid W))\phi_{j}(W):j\}$. In $L^{2}(P)$, functions of $W$ are orthogonal to the linear span of $\{(A-g(1\mid W))\phi_{j}(W):j\}$. Therefore, the projection of $\bar{Q}_{P}$ on this orthogonal sum space is given by $\theta_{P}(W)=E(\bar{Q}_{P}\mid W)=E_{P}(Y\mid W)$ plus the projection of $\bar{Q}_{P}=\tilde{\theta}_{P}+A\tau_{A,P}$ onto $\{(A-g(1\mid W))\phi_{j}(W):j\}$. The projection of $\tilde{\theta}_{P}$ on this linear span equals zero since a function of $W$ is orthogonal to $(A-g(1\mid W))\phi_{j}(W)$. The projection of $A\tau_{A,P}$ is given by $(A-g(1\mid W))\sum_{j}\beta_{P}(j)\phi_{j}$ with $\beta_{P}=\arg\min_{\beta}E_{P}(A\tau_{A,P}(W)-(A-g(1\mid W))\sum_{j}\beta(j)\phi_{j}(W))^{2}$. We can write $A\tau_{A,P}=(A-g(1\mid W))\tau_{A,P}(W)+g(1\mid W)\tau_{A,P}(W)$. So $\beta_{P}=\arg\min_{\beta}E_{P}(A-g(1\mid W))^{2}(\tau_{A,P}(W)-\sum_{j}\beta(j)\phi_{j}(W))^{2}$, which equals $\arg\min_{\beta}E_{P}g_{P}(1-g_{P})(1\mid W)(\tau_{A,P}(W)-\sum_{j}\beta(j)\phi_{j}(W))^{2}$. This proves the lemma. ∎
@@
We want to derive the canonical gradient of $\beta_{P}$ defined on a nonparametric model. Our starting point is that
$\beta_{P}=\arg\min_{\beta}E_{P}\left(Y-(A-g_{P}(1\mid W))\sum_{j}\beta(j)\phi_{j}(W)\right)^{2}.$
@@
$U(P,g_{P},\beta)=E_{P}(A-g_{P}(1\mid W)){\bf\phi}(W)\left(Y-(A-g_{P}(1\mid W))\sum_{j}\beta(j)\phi_{j}(W)\right)=0.$
@@
where $D_{1,P}$ is the canonical gradient of $P\rightarrow U(\beta_{1},P,g_{1})$ at $\beta_{1}=\beta_{P}$ and $g_{1}=g_{P}$, and $D_{2,P}$ is the canonical gradient of $P\rightarrow U(\beta_{1},P_{1},g_{P})$ at $\beta_{1}=\beta_{P}$ and $P_{1}=P$. Since $U(\beta_{1},P,g_{1})$ is just a mean parameter with respect to $P$ (like $E_{P}D$ which has canonical gradient $D-E_{P}D$), it follows that
$D_{1,P}=(A-g_{P}(1\mid W)){\bf\phi}(W)\left(Y-(A-g_{P}(1\mid W))\sum_{j}\beta_{P}(j)\phi_{j}(W)\right).$
@@
$U(\beta_{1},P_{1},g_{P})=E_{P_{1}}(A-g_{P}(1\mid W)){\bf\phi}(W)\left(\bar{Q}_{P_{1}}-(A-g_{P}(1\mid W))\sum_{j}\beta_{P_{1}}(j)\phi_{j}\right).$
@@
$=\int_{w,a}p_{1}(w,a)(a-g_{P}(1\mid w)){\bf\phi}(w)\bar{Q}_{1}(w,a)-\int_{w,a}p_{1}(w,a)(a-g_{P}(1\mid w))^{2}{\bf\phi}(w)\sum_{j}\beta_{1}(j)\phi_{j}(w)$
at $p_{1}=p$. We know that the canonical gradient of $P\rightarrow g_{P}(1\mid w)$ is given by $I(W=w)/p(w)(A-g_{P}(1\mid W))$. So, by the delta-method, the canonical gradient is given by
@@
$D_{1,P}+D_{2,P}$ $=(A-g_{P}(1\mid W)){\bf\phi}(W)\left(Y-(A-g_{P}(1\mid W))\sum_{j}\beta_{P}(j)\phi_{j}(W)\right)-(A-g_{P}(1\mid W)){\bf\phi}(W)\theta_{P}(W)$
$=(A-g_{P}(1\mid W)){\bf\phi}(W)\left(Y-\theta_{P}(W)-(A-g_{P}(1\mid W))\sum_{j}\beta_{P}(j)\phi_{j}(W)\right).$